# Densing Law of LLMs

### v1

@@ Abstract
Large Language Models (LLMs) have emerged as a milestone in artificial intelligence, and their performance can improve as the model size increases. However, this scaling brings great challenges to training and inference efficiency, particularly for deploying LLMs in resource-constrained environments, and the scaling trend is becoming increasingly unsustainable. This paper introduces the concept of “capacity density” as a new metric to evaluate the quality of the LLMs across different scales and describes the trend of LLMs in terms of both effectiveness and efficiency. To calculate the capacity density of a given target LLM, we first introduce a set of reference models and develop a Scaling Law to predict the downstream performance of these reference models based on their parameter sizes. We then define the effective parameter size of the target LLM as the parameter size required by a reference model to achieve equivalent performance, and formalize the capacity density as the ratio of the effective parameter size to the actual parameter size of the target LLM. Capacity density provides a unified framework for assessing both model effectiveness and efficiency. Our further analysis of recent open-source base LLMs reveals an empirical law (the Densing Law) that the capacity density of LLMs grows exponentially over time. More specifically, using some widely used benchmarks for evaluation, the capacity density of LLMs doubles approximately every three months. The law provides new perspectives to guide future LLM development, emphasizing the importance of improving capacity density to achieve optimal results with minimal computational overhead.
@@
We introduce the concept of “capacity density” to evaluate the training quality of large language models (LLMs) and describe the trend of LLMs that considers both effectiveness and efficiency.
(Relative) Capacity Density.
For a given LLM $\mathcal{M}$, its capacity density is defined as the ratio of its effective parameter size to its actual parameter size, where the effective parameter size is the minimum number of parameters required for the reference model to achieve performance equivalent to $\mathcal{M}$.
We reveal an empirical law for the capacity density of open-source base LLMs released since 2023.
@@
The maximum capacity density of LLMs exhibits an exponential growth trend over time. $\text{ln}(\rho_{\text{max}})=At+B$ Here, $\rho_{\text{max}}$ is the maximum capacity density of LLMs at time $t$.
Figure 1 presents the capacity density of popular LLMs, measured by their performance on $5$ widely-used benchmarks. A trend is fitted between maximum capacity density and release date, revealing that $A\approx 0.007$ with $R^{2}\approx 0.93$. This indicates the maximum capacity density of LLMs doubles approximately every 3.3 months [^1] The capacity density growth rate is affected by specific evaluation benchmarks and reference models.. In other words, around three months, it is possible to achieve performance comparable to current state-of-the-art LLMs using a model with half the parameter size.
Figure 1: The estimated capacity density of open-source base LLMs.
@@ 1 Introduction
To this end, we introduce the concept of capacity density, which serves as a metric for evaluating and comparing the training quality of LLMs on various scales. Accurately measuring all aspects of an LLM’s capabilities, or its level of intelligence, is quite challenging. In this article, we design a method to assess the relative capacity density [^2] For ease of explanation, in this work, we use “density” to refer to “(relative) capacity density”.. Specifically, we use a reference model and then estimate its scaling function between the performance on downstream tasks and parameter sizes. Based on the scaling function, for any given model, we calculate its effective parameter size – the number of parameters the reference model would need to achieve equivalent performance. The density of an LLM relative to the reference model is then defined as the ratio of its effective parameter size to its actual parameter size. By introducing the concept of model density, we aim to more accurately measure model quality and enable comparisons between models of different scales. This evaluation method has the potential to provide new insights into the future direction of LLM development, helping researchers find the optimal balance between effectiveness and efficiency.
@@ 1.1 Key Findings
The maximum capacity density of LLMs exhibits an exponential growth trend over time. $\text{ln}(\rho_{\text{max}})=A\cdot t+B$ Here, $\rho_{\text{max}}$ is the maximum capacity density of LLMs at time $t$.
Based on our evaluation on $5$ widely-used benchmarks, MMLU (Hendrycks et al., 2020), BBH (Suzgun et al., 2023), MATH (Hendrycks et al., 2021), HumanEval (Chen et al., 2021), and MBPP (Austin et al., 2021), $A\approx 0.007$, which means the maximum density of LLMs doubles approximately every three months. For example, MiniCPM-1-2.4B released on February 1st, 2024, can achieve comparable or even superior performance with Mistral-7B released on September 27th, 2023. We can use an LLM with only 3535% parameters to obtain roughly equivalent performance after $4$ months. It is worth noting that using different evaluation benchmarks may result in slight variations in the estimation and growth rate of model density. We encourage the community to develop more comprehensive evaluation benchmarks for LLMs to ensure more accurate measurements of density.
@@
In this work, we propose a new evaluation metric, density, for LLMs, which can offer a new, unified perspective on the two current trends – enhancing effectiveness and increasing efficiency. Based on our proposed metric, we evaluate $29$ open-source models and find an empirical experience law, named Densing Law: the density of LLMs exhibits an exponentially increasing trend. Based on this empirical relationship, we discuss several deductions and provide observational evidence. Through this novel evaluation perspective, we hope to provide valuable insights and guidance for the future development of LLMs.
@@ 2.4 Density
$\hat{\mathcal{L}}(S_{\mathcal{M}})=l-\frac{1}{\gamma}ln\left(\frac{c}{S_{\mathcal{M}}-d}-1\right);\quad\hat{N}(S_{\mathcal{M}})=\left(\frac{\hat{\mathcal{L}}(S_{\mathcal{M}})-bD_{0}^{-\beta}}{a}\right)^{-\frac{1}{\alpha}}.$ (4)
@@ 3.1 Evaluation Settings
Dataset In this work, we adopt the following widely-used datasets for evaluation: MMLU (Hendrycks et al., 2020) for English knowledge-intensive tasks, BBH (Suzgun et al., 2023) for challenging logic reasoning tasks, MATH (Hendrycks et al., 2021) for mathematical reasoning tasks, and HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021) for coding tasks. We apply the open-source tools (OpenCompass, 2023; Liu et al., 2024) for evaluation. Here, we evaluate all models in a few-shot in-context learning manner and these models are required to generate the final answer label based on the given demonstrations and inputs of test instances. Following widely-used settings, MMLU, BBH, MATH, HumanEval, and MBPP are evaluated under the $5$-shot, 33-shot, 44-shot, 0-shot, and $3$-shot settings, respectively. Besides, for BBH, MATH, and MBPP, we adopt the chain-of-thought prompting technique (Wei et al., 2022b).
@@ 4 Discussion
Accurate Capacity Measurement Capability density reflects the abilities of an LLM per unit of parameters. However, with current technology, we cannot accurately assess the absolute capability level of LLMs, meaning that quantifying intelligence remains a great challenge. Therefore, in this work, we design a method to measure the relative density value of LLMs. Besides, we use widely-used benchmarks to evaluate the performance of LLMs. However, the limited number of benchmarks and potential data contamination issues introduce bias in performance evaluation. Thus, advancing accurate measurement of LLMs’ capabilities or intelligence levels in the future will enable better calculation of their density.
@@ 5 Limitations and Future Directions
In this section, we discuss the limitations and future directions of our proposed method to evaluate the capacity density of LLMs.
Fair and Comprehensive Evaluation The capacity density measurement of LLMs relies on existing benchmarks to evaluate model performance. Therefore, the benchmark quality greatly impacts the density measurement results. In this work, we use those benchmarks widely adopted by researchers to evaluate various LLMs. However, several challenges remain: (1) Comprehensive evaluation: With the development of LLMs, the capabilities of LLMs significantly expand, such as the ability to handle complex reasoning tasks (OpenAI, 2024b). Consequently, the capacity density measurement needs to be continually updated by incorporating more comprehensive evaluation datasets that reflect evolving capabilities. (2) Fair evaluation: With the increasing scale of pre-training data and the construction of synthetic data, some LLMs are overoptimized towards benchmarks, leading to inflated scores. To address this, we plan to use newly constructed datasets to evaluate model performance, thereby mitigating the overfitting risk and ensuring accurate density estimation.
Multi-modal Density In this work, we focus on measuring the capacity density of language models. However, measuring the density and trends in large multimodal models is also crucial as multimodal applications increase. In the future, designing reasonable density evaluation methods for multimodal models will be an important research direction.
Inference Densing Law Recent research has highlighted that more inference computational costs allow LLMs to engage in deeper reasoning, effectively enhancing their performance on complex tasks (OpenAI, 2024b). In this work, we use the parameter size as the basis to evaluate model capacity density. However, as the importance of chain-of-thought reasoning continues to grow, density evaluation should shift towards being based on inference FLOPs. Specifically, capacity density could be formalized as the ratio of effective inference FLOPs to actual inference FLOPs. In this way, we hope that LLMs achieve optimal results with the minimum number of reasoning steps.
@@ 6 Conclusion
To illustrate the recent trend towards efficient LLMs and to quantitatively measure the training quality of LLMs, this paper introduces a method for evaluating the capacity density of LLMs. By measuring the capacity density of open-source base LLMs released since 2023, we show an empirical law: the capacity density of LLMs increases exponentially over time. The evaluation results on some widely-used LLM benchmarks indicate that the density of LLMs doubles every three months. This implies that, within three months, a model with only half the parameters can achieve performance comparable to the current state-of-the-art models. This finding highlights the rapid development and increasing efficiency of LLMs. We discuss several corollaries based on the law, and hope that the law and its corollaries will encourage the LLM community to continue enhancing model capacity density and achieving optimal performance with minimal computational costs.

### v2

@@ Abstract
Large Language Models (LLMs) have emerged as a milestone in artificial intelligence, and their performance can improve as the model size increases. However, this scaling brings great challenges to training and inference efficiency, particularly for deploying LLMs in resource-constrained environments, and the scaling trend is becoming increasingly unsustainable. This paper introduces the concept of “capability density” as a new metric to evaluate the quality of the LLMs across different scales and describes the trend of LLMs in terms of both effectiveness and efficiency. To calculate the capability density of a given target LLM, we first introduce a set of reference models and develop a Scaling Law to predict the downstream performance of these reference models based on their parameter sizes. We then define the effective parameter size of the target LLM as the parameter size required by a reference model to achieve equivalent performance, and formalize the capability density as the ratio of the effective parameter size to the actual parameter size of the target LLM. Capability density provides a unified framework for assessing both model effectiveness and efficiency. Our further analysis of recent open-source base LLMs reveals an empirical law (Densing Law) that the capability density of LLMs grows exponentially over time. More specifically, using some widely used benchmarks for evaluation, the capability density of LLMs doubles approximately every three months. The law provides new perspectives to guide future LLM development, emphasizing the importance of improving capability density to achieve optimal results with minimal computational overhead.
@@
We introduce the concept of “capability density” to evaluate the training quality of large language models (LLMs) and describe the trend of LLMs that considers both effectiveness and efficiency.
(Relative) Capability Density.
For a given LLM $\mathcal{M}$, its capability density is defined as the ratio of its effective parameter size to its actual parameter size, where the effective parameter size is the minimum number of parameters required for the reference model to achieve performance equivalent to $\mathcal{M}$.
We reveal an empirical law for the capability density of open-source base LLMs released since 2023.
@@
The maximum capability density of LLMs exhibits an exponential growth trend over time. $\text{ln}(\rho_{\text{max}})=At+B$ Here, $\rho_{\text{max}}$ is the maximum capability density of LLMs at time $t$.
Figure 1 presents the capability density of popular LLMs, measured by their performance on $5$ widely-used benchmarks. A trend is fitted between maximum capability density and release date, revealing that $A\approx 0.007$ with $R^{2}\approx 0.93$. This indicates the maximum capability density of LLMs doubles approximately every 3.3 months [^1] The capability density growth rate is affected by specific evaluation benchmarks and reference models.. That means, around three months, it is possible to achieve performance comparable to current state-of-the-art LLMs using a model with half the parameter size.
Figure 1: The estimated capability density of open-source base LLMs.
@@ 1 Introduction
To this end, we introduce the concept of capability density, which serves as a metric for evaluating and comparing the training quality of LLMs on various scales. Accurately measuring all aspects of an LLM’s capabilities, or its level of intelligence, is quite challenging. In this article, we design a method to assess the relative capability density [^2] For ease of explanation, in this work, we use “density” to refer to “(relative) capability density”.. Specifically, we use a reference model and then estimate its scaling function between the performance on downstream tasks and parameter sizes. Based on the scaling function, for any given model, we calculate its effective parameter size – the number of parameters the reference model would need to achieve equivalent performance. The density of an LLM relative to the reference model is then defined as the ratio of its effective parameter size to its actual parameter size. By introducing the concept of model density, we aim to more accurately measure model quality and enable comparisons between models of different scales. This evaluation method has the potential to provide new insights into the future direction of LLM development, helping researchers find the optimal balance between effectiveness and efficiency.
@@ 1.1 Key Findings
The maximum capability density of LLMs exhibits an exponential growth trend over time. $\text{ln}(\rho_{\text{max}})=A\cdot t+B$ Here, $\rho_{\text{max}}$ is the maximum capability density of LLMs at time $t$.
Based on our evaluation on $5$ widely-used benchmarks, MMLU (Hendrycks et al., 2020), BBH (Suzgun et al., 2023), MATH (Hendrycks et al., 2021), HumanEval (Chen et al., 2021), and MBPP (Austin et al., 2021), $A\approx 0.007$, which means the maximum density of LLMs doubles approximately every three months. For example, MiniCPM-1-2.4B released on February 1st, 2024, can achieve comparable or even superior performance with Mistral-7B released on September 27th, 2023. We can use an LLM with only $35$% parameters to obtain roughly equivalent performance after $4$ months. It is worth noting that using different evaluation benchmarks may result in slight variations in the estimation and growth rate of model density. We encourage the community to develop more comprehensive evaluation benchmarks for LLMs to ensure more accurate measurements of density.
@@
In this work, we propose a new evaluation metric, capability density, for LLMs, which can offer a new, unified perspective on the two current trends – enhancing effectiveness and increasing efficiency. Based on our proposed metric, we evaluate $29$ open-source models and find an empirical experience law, named Densing Law: the density of LLMs exhibits an exponentially increasing trend. Based on this empirical relationship, we discuss several deductions and provide observational evidence. Through this novel evaluation perspective, we hope to provide valuable insights and guidance for the future development of LLMs.
@@ 2.4 Density
^$\hat{\mathcal{L}}(S_{\mathcal{M}})=l-\frac{1}{\gamma}ln\left(\frac{c}{S_{\mathcal{M}}-d}-1\right);\quad\hat{N}(S_{\mathcal{M}})=\left(\frac{\hat{\mathcal{L}}(S_{\mathcal{M}})-bD_{0}^{-\beta}}{a}\right)^{-\frac{1}{\alpha}}.$ (4)
@@ 3.1 Evaluation Settings
Dataset In this work, we adopt the following widely-used datasets for evaluation: MMLU (Hendrycks et al., 2020) for English knowledge-intensive tasks, BBH (Suzgun et al., 2023) for challenging logic reasoning tasks, MATH (Hendrycks et al., 2021) for mathematical reasoning tasks, and HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021) for coding tasks. We apply the open-source tools (OpenCompass, 2023; Liu et al., 2024) for evaluation. Here, we evaluate all models in a few-shot in-context learning manner and these models are required to generate the final answer label based on the given demonstrations and inputs of test instances. Following widely-used settings, MMLU, BBH, MATH, HumanEval, and MBPP are evaluated under the $5$-shot, $3$-shot, $4$-shot, 00-shot, and $3$-shot settings, respectively. Besides, for BBH, MATH, and MBPP, we adopt the chain-of-thought prompting technique (Wei et al., 2022b).
@@ 4 Discussion
Accurate Capability Measurement Capability density reflects the abilities of an LLM per unit of parameters. However, with current technology, we cannot accurately assess the absolute capability level of LLMs, meaning that quantifying intelligence remains a great challenge. Therefore, in this work, we design a method to measure the relative density value of LLMs. Besides, we use widely-used benchmarks to evaluate the performance of LLMs. However, the limited number of benchmarks and potential data contamination issues introduce bias in performance evaluation. Thus, advancing accurate measurement of LLMs’ capabilities or intelligence levels in the future will enable better calculation of their density.
@@ 5 Limitations and Future Directions
In this section, we discuss the limitations and future directions of our proposed method to evaluate the capability density of LLMs.
Fair and Comprehensive Evaluation The capability density measurement of LLMs relies on existing benchmarks to evaluate model performance. Therefore, the benchmark quality greatly impacts the density measurement results. In this work, we use those benchmarks widely adopted by researchers to evaluate various LLMs. However, several challenges remain: (1) Comprehensive evaluation: With the development of LLMs, the capabilities of LLMs significantly expand, such as the ability to handle complex reasoning tasks (OpenAI, 2024b). Consequently, the capability density measurement needs to be continually updated by incorporating more comprehensive evaluation datasets that reflect evolving capabilities. (2) Fair evaluation: With the increasing scale of pre-training data and the construction of synthetic data, some LLMs are overoptimized towards benchmarks, leading to inflated scores. To address this, we plan to use newly constructed datasets to evaluate model performance, thereby mitigating the overfitting risk and ensuring accurate density estimation.
Multi-modal Density In this work, we focus on measuring the capability density of language models. However, measuring the density and trends in large multimodal models is also crucial as multimodal applications increase. In the future, designing reasonable density evaluation methods for multimodal models will be an important research direction.
Inference Densing Law Recent research has highlighted that more inference computational costs allow LLMs to engage in deeper reasoning, effectively enhancing their performance on complex tasks (OpenAI, 2024b). In this work, we use the parameter size as the basis to evaluate model capability density. However, as the importance of chain-of-thought reasoning continues to grow, density evaluation should shift towards being based on inference FLOPs. Specifically, capability density could be formalized as the ratio of effective inference FLOPs to actual inference FLOPs. In this way, we hope that LLMs achieve optimal results with the minimum number of reasoning steps.
@@ 6 Conclusion
To illustrate the recent trend towards efficient LLMs and to quantitatively measure the training quality of LLMs, this paper introduces a method for evaluating the capability density of LLMs. By measuring the capability density of open-source base LLMs released since 2023, we show an empirical law: the capability density of LLMs increases exponentially over time. The evaluation results on some widely-used LLM benchmarks indicate that the density of LLMs doubles every three months. This implies that, within three months, a model with only half the parameters can achieve performance comparable to the current state-of-the-art models. This finding highlights the rapid development and increasing efficiency of LLMs. We discuss several corollaries based on the law, and hope that the law and its corollaries will encourage the LLM community to continue enhancing model capability density and achieving optimal performance with minimal computational costs.